OpenStudy (anonymous):

Part 1: Solve |4x-11| < 5. You must show all work. (3 points) Part 2: Use complete sentences to describe the graph in words.

OpenStudy (anonymous):

thats all it says

hero (hero):

Hint: In general: \(|x - a| < b \equiv -b < x - a < b \equiv a - b < x < a + b \)

OpenStudy (anonymous):

so whats the answer can you explain it too me

hero (hero):

\(\equiv\) means "equivalent to"

OpenStudy (anonymous):

i mean step by step process

hero (hero):

Written another way: \(|x - a| < b \\ = -b < x - a < b \\ = a - b < x < a + b\)

OpenStudy (anonymous):

so whats the answer

hero (hero):

You should come here looking for explanations, not answers.

OpenStudy (anonymous):

but you not explaining throughly like walk me throught it

hero (hero):

Hint: a = 11, b = 5

OpenStudy (anonymous):

|4x-11|<5 4x<16 x<4 or |4x-11|>-5 4x>6 x>2/3

hero (hero):

You changed it from an "and" statement to an "or" statement. But it's not an "or" statement.

OpenStudy (anonymous):

so its |4x-11|<5 4x<16 x<4 and |4x-11|>-5 4x>6 x>2/3

hero (hero):

|4x-11| < 5 -5 < 4x - 11 < 5 11 - 5 < 4x - 11 < 11 + 5 6 < 4x < 16 6/4 < x < 16/4 2/3 < x < 4 That's how it should be done.